- Status Report
- Jan 31, 2023
ISS On-orbit Status 18 Jan 2003
All ISS systems continue to function nominally, except as noted previously
or below. It is another Saturday, the first of two crew off-duty days.
After wake-up at 1:00am EST, morning inspection, morning hygiene and
breakfast, the station residents completed the weekly 3-hr. house cleaning.
[The "uborka stantsii" focuses on removal of food waste products,
cleaning of compartments with vacuum cleaner, wet cleaning of surfaces
with disinfectants and cleaning of fan screens to avoid temperature rises.]
FE-1 Nikolai Budarin performed the daily routine inspection of the Rasteniya-2/Lada-2
(BIO-5) plant growth experiment.
FE-2/SO Don Pettit took carbon dioxide (CO2) readings in the Service
Module (SM) and Lab with the CDM (CO2 monitor). [This is in support of
the current measurement program to help resolve conflicting ppCO2 (carbon
dioxide partial pressure) readings by the SM gas analyzer and the U.S.
MCA (major constituents analyzer), which yesterday underwent another
full ("zero"-) calibration.]
At the BMP micropurification system in the SM, Nikolai Budarin terminated
regeneration mode for absorption bed #1, switching it back to Purify
mode. Later, he initiated the regenerative bake-out of catalyst bed #2.
Later, Nikolai gave the SM’s RLS amateur radio equipment, with its VHF
2m transceiver and Packet module, a quick performance checkout. At 11:55am,
CDR Ken Bowersox set up the equipment for a ham radio pass with high
school students at Cape Cod National Seashore School, Wellfleet, Mass.,
which was conducted at 12:00pm. [The U.S. Park Service has developed
partnerships with local amateur radio groups to draw attention to the
cultural and scientific values of the Marconi Station Site as well as
the current 100th Anniversary of Marconi’s trans-Atlantic message sent
from the U.S. (Cape Cod) to Europe in 1903, demonstrating the first practical
system of wireless telegraphy. Thousands of students, community groups,
and park visitors have participated in Cape Cod National Seashore’s "Marconi" activities
to prepare for an "on the air" worldwide radio party currently
taking place (Jan 11-19). Guglielmo Marconi’s daughter, Princess Elettra
Marconi joined the students, plus representatives from the Italian Embassy.]
Regular daily maintenance of the SOSh environment control and life support
system and the IMS inventory update auto import/export preparation were
completed by Pettit, while Bowersox did the autonomous payload status
check in the Lab.
Bowersox also performed the planned deactivation of the ZCG (Zeolite
crystal growth) furnace heaters in the Lab.
The crew worked out on RED (resistive exercise device), TVIS (treadmill
with vibration isolation and stabilization), CEVIS (cycle ergometer with
vibration isolation) and VELO (Russian cycle ergometer) with NS load
trainer, performing the regular daily 2.5-hr. exercise.
During the day, Huntsville POC (Payload Operations Center) planned to
command a power-up of EXPRESS Rack 5 (ER5) of about 4 hrs. to determine
the status of the new software loaded onto the rack yesterday, check
POC capability to configure ER5 via ground commanding and verify the
health and science telemetry from the rack. [The crew was thanked for
their good work during the RIC (rack interface controller) software upload,
hoping that problems associated with the software uploads have all been
Yesterday, EATCS (external active thermal control system) loop B heaters
1 and 2 for the NTA (nitrogen tank assembly) experienced a failure whose
root cause is currentlyunknown. Lower temperature limits of NTA components
hardware were not violated at any time. Troubleshooting from the ground
is underway. [Up to the time of the failure, heater 1 had nominally cycled
between its lower and upper setpoints, providing the thermal conditioning
of the loop B NTA. At 11:07am EST, the heater 1 RPC (remote power controller)
closed as expected, as temperature reached its lower setpoint, but it
did not start to increase. This resulted in an "undertemp" caution
annunciation. Subsequently, the heater 2 RPC closed as temperature reachedthat
heater’slower setpoint. Again, the temperatures did not increase, which
also resulted in an "undertemp" caution.]
When the MSG (microgravity science glovebox) repair parts PDC and ESEM3
have reached Baikonur on 1/24-25, they will be accommodated for late
load in Progress M-247 (10P). This means, they may get stowed in the
side hatch and may not immediately be accessible on unloading. [After
MSG reactivation, first experiment to be processed will be the In-SPACE
payload. Plans also call for one CSLM sample to be processed (the CSLM
run will be conducted to verify MSG homogenous temperature distribution
andmaintenance of the vacuum in the sample). Spillover of MSG science
into Increment 7 will be limited to PFMI (pore formation and migration
Science Update (Expedition Six — 7th):
The crew was thanked for conducting science operations this week, viz.,
CEO, MISSE photos during the EVA, PuFF, ZCG and upgrading the EXPRESS
Rack 5 software..
Extra-Vehicular Activity Radiation Monitors (EVARM): n/a
GASMAP/Pulmonary Function in Flight (PuFF): The GASMAP 90 Day Health
Check on 1/21 will include an auto calibration in place of the regular
Renal (Kidney) Stone Experiment: Continuing nominally (crew taking pills).
Human Research Facility/Workstation (HRF WS): Continuing.
Foot/Ground Reaction Forces During Space Flight (FOOT): The FOOT team
will be standing by if Bowersox needs a hand during his next session.
Pore Formation and Mobility Investigation (PFMI): Complete for Inc.
5. On hold pending MSG PDC (power distribution controller) and ESEM3
(exchangeable standard electronic module 3) replacement. Will remain
in MSG until further notice.
Space Acceleration Measurement System (SAMS): SAMS captured the EVA
on 1/15 and continues to collect acceleration data for vibratory characterization
of the microgravity environment.
Microgravity Acceleration Measurement System (MAMS): MAMS captured the
EVA on 1/15 and continues measurement of microgravity environment in
the quasi-steady regime for general characterization.
Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):
Temperatures are nominal. PCG-STES is currently in heating mode.
Materials ISS Experiment (MISSE): In progress. Deployed outside. Nominal
and collecting data. Was photographed by the crew during the 1/15 EVA.
Zeolite Crystal Growth (ZCG): ZCG is operating nominally. The ground
is looking forward to getting samples back from both runs for analysis.
Thank you to the crew for all the hard work they put into completing
EarthKAM (EK): Anxiously awaiting resumption of activities on 1/28.
Crew Earth Observation (CEO): Continuing to receive and review downlinked
ECS (electronic still camera) imagery over southern hemisphere targets.
Many thanks to the crew’s ongoing efforts and dedication to acquire these
Today’s CEO targets were Congo-Zimbabwe Biomass Burning (although this
is the low-fire [wet] season, some fires may exist, and burn scars from
previous seasons are of interest [as irregular shapes often superimposed,
so that the most recent fires are the clearest, several fire seasons
being recorded in fire-prone areas]. Looking mainly right for oblique
views), Lower Amazon River Basin (nadir pass over the Amazon River delta
region. Clearing weather at the basin and inland should have yielded
nice opportunity to document dynamic estuary changes in river just right
of ISS track), Perth, Australia (Astronaut John Glenn called Perth the "City
of Lights" after his historic fly-over in 1962. High pressure will
dominate the area off the coast of Western Australia allowing for excellent
viewing opportunity of Perth suburbs and the port city of Fremantle just
south of track), and Patagonian Glaciers (nice pass between two slow
moving frontal systems to capture detailed views of glacier shape and
surface morphology [crevasse fields appear as aligned cracks], and color.
Views of glacier tongues and peaks will all be north of ISS track.)
CEO images can be viewed at the website
ISS Orbit (as of this morning, 7:28am EST [= epoch]):
Mean altitude — 389.6 km
Apogee — 393.5 km
Perigee — 385.7 km
Period — 92.35 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005792
Orbits per 24-hr. day — 15.59
Altitude loss — 170 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 23765
For more on ISS orbit and worldwide naked-eye visibility dates/times,